# What Is That Talk About? A Video-to-Text Summarization Dataset for Scientific Presentations

### v1

@@ 3 VISTA Dataset
Table 1 compares VISTA with several existing video-to-text summarization datasets. While many focus on open-domain (e.g., MMSum, Instruct-V2Xum) or focus on specific areas like news (MLASK, MM-AVS) and activities (VideoXum), VISTA is tailored for summarizing scientific presentations, addressing a distinct niche in video-to-text summarization. On average, it features longer inputs (6.8 minutes) than VideoXum (2.1 minutes) and BLiSS (5.0 minutes), as well as longer summaries (192.6 tokens), compared to YouCook2 (67.8 tokens) and VideoXum (49.9 tokens).
Table 2 summarizes the dataset statistics: videos average 6.76 minutes and 16.36 shots (we use PySceneDetect with ContentDetector to calculate video shots), while summaries contain 192.62 tokens on average across 7.19 sentences. The average dependency tree depth (Avg. Depth of Dep Tree) is 6.02, indicating the syntactic complexity of the summaries, while the Type-Token Ratio (TTR) of 0.62 reflects lexical diversity, calculated using spaCy. Diversity metrics Li et al. (2016), which measure the variety of unique n-grams, yield Distinct-1, Distinct-2, and Distinct-3 scores of 0.62, 0.93, and 0.97, respectively. Figure 3 visualizes key attributes: most summaries remain under 250 tokens and 10 sentences, and most videos last fewer than 10 minutes with under 30 shots. In Appendix C, we present two random samples from the VISTA dataset.
@@ 4 Benchmarking VISTA
A key challenge in video-to-text summarization is structuring the generated summaries in a coherent and faithful manner. Directly learning the mapping from $v$ to $s$ could lead to inadequate outputs, as the model lacks explicit guidance on how to organize and present the extracted information. Scientific abstracts often follow a relatively well-defined structure, making them suitable for a more structured generation approach Takeshita et al. (2024). We follow previous work Narayan et al. (2021); Liu et al. (2023); Narayan et al. (2023) in adopting a plan-based framework that introduces an intermediate representation to capture latent structure more effectively than simpler end-to-end approaches. Specifically, given input video $v$, we first generate plan $p$, which consists of a sequence of automatically generated questions $\{q_{1},q_{2},\ldots,q_{m}\}$, each corresponding to a sentence to be verbalized in the summary. The plan explicitly controls the structure of the summary as a whole and the content of each of its sentences (which are meant to answer the questions in the plan). The model is then trained to learn the extended conditional probability distribution $P(s\leavevmode\nobreak\ |\leavevmode\nobreak\ v,p)$, ensuring that the generated summaries follow the structure and flow of plan $p$.
@@
We hypothesize that summary sentences can be viewed as responses to plan questions directly associated with them. This idea is inspired by the theory of Questions Under Discussion Roberts (2012); Wu et al. (2023b); Suvarna et al. (2024), which posits that discourse often revolves around a set of questions that guide the structure and interpretation of the conversation.
@@ 6 Results and Analysis
We analyze the plan generation ablation strategy by comparing it with simpler baselines: Lead-3Q, Tail-3Q, and Random-3Q. In these ablation baselines, plans are generated by selecting the first three, last three, or three randomly chosen summary sentences, respectively. Each selected sentence serves as a target for generating a question, with its preceding sentences providing the context. For instance, in the Lead-3Q setting, the first sentence is used as the target (without any preceding context), prompting the first question in the plan, while subsequent sentences incorporate earlier ones as context.
@@
Table 4 underlines the performance differences across different plan generation ablation strategies. The Lead-3Q strategy performs better overall compared to Tail-3Q and Random-3Q, indicating that initial sentences offer stronger contextual continuity for generating plan questions. Nonetheless, these heuristic strategies fail to match the performance of the original planning method.
@@ 8 Conclusion
This paper introduces VISTA, a dataset for summarizing scientific video presentations into concise textual summaries. Comprehensive evaluations across multiple large models demonstrate that the summarization task is challenging, relying on the interplay of multiple modalities (video, text, and audio). We further introduce a plan-based approach, which yields improvements in summary quality and factual accuracy. Beyond dataset creation, our work also confirms that current leading large models still exhibit a noticeable gap compared to human performance. In the future, we would like to explore more specialized models for analyzing tables, charts, and figures. We also plan to investigate whether summarizing the video chunk-by-chunk could lead to more faithful output.
@@
We use several state-of-the-art large models in our experiments and select the best-performing model, mPLUG-Owl3, to demonstrate the effectiveness of the planning strategy. These large models may carry biases introduced during pretraining. We have not assessed the extent of these biases, as they lie beyond the scope of this study. Furthermore, we have not tested the plan-based approach on all model variants presented in our experiments (e.g., text-based large models and audio-based large models). Our work does not aim to prove that the plan-based method is effective in all models of different modalities, but rather to demonstrate that the plan-based method can improve the performance of video-based models on our dataset. Future work could examine how the plan-based method performs across a wider range of models and modalities.
@@ Acknowledgements
This project has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 Research and Innovation Programme (Grant Agreement No. 948878).
@@ Appendix C Data Sample
The VISTA dataset contains carefully curated video-text pairs, predominantly sourced from published papers, aiming to ensure a high standard of quality and relevance. The accompanying texts are designed to function as summaries of their respective videos, offering a concise representation of their content (see Figure 7 and Figure 8). Additionally, our dataset focuses on topics within the field of artificial intelligence, making it a good resource for research in AI-related video-to-text summarization and comprehension.
@@
Refer to caption An emerging solution for explaining Transformer-based models is to use vector-based analysis on how the representations are formed. However, providing a faithful vector-based explanation for a multi-layer model could be challenging in three aspects: (1) Incorporating all components into the analysis, (2) Aggregating the layer dynamics to determine the information flow and mixture throughout the entire model, and (3) Identifying the connection between the vector-based analysis and the model’s predictions. In this paper, we present DecompX to tackle these challenges. DecompX is based on the construction of decomposed token representations and their successive propagation throughout the model without mixing them in between layers. Additionally, our proposal provides multiple advantages over existing solutions for its inclusion of all encoder components (especially nonlinear feed-forward networks) and the classification head. The former allows acquiring precise vectors while the latter transforms the decomposition into meaningful prediction-based values, eliminating the need for norm- or summation-based vector aggregation. According to the standard faithfulness evaluations, DecompX consistently outperforms existing gradient-based and vector-based approaches on various datasets.
Figure 8: A random sample from the VISTA dataset, originating from Modarressi et al. (2023).

### v2

@@ 3 VISTA Dataset
Table 1 compares VISTA with several existing video-to-text summarization datasets. While many focus on open-domain (e.g., MMSum, Instruct-V2Xum) or focus on specific areas like news (MLASK, MM-AVS) and activities (VideoXum), VISTA is tailored for summarizing scientific presentations, addressing a distinct niche in video-to-text summarization. On average, it features longer inputs (6.8 minutes) than VideoXum (2.1 minutes) and MSS (3.4 minutes), as well as longer summaries (192.6 tokens), compared to YouCook2 (67.8 tokens) and VideoXum (49.9 tokens).
Table 2 summarizes the dataset statistics: videos average 6.76 minutes and 16.36 shots (we use PySceneDetect with ContentDetector to calculate video shots), while summaries contain 192.62 tokens on average across 7.19 sentences. The average dependency tree depth (Avg. Depth of Dep Tree) is 6.02, indicating the syntactic complexity of the summaries. Meanwhile, the Type-Token Ratio (TTR) is 0.62, reflecting lexical diversity. Both metrics are calculated using spaCy. Diversity metrics Li et al. (2016), which measure the variety of unique n-grams, yield Distinct-1, Distinct-2, and Distinct-3 scores of 0.62, 0.93, and 0.97, respectively. Figure 3 visualizes key attributes: most summaries remain under 250 tokens and 10 sentences, and most videos last fewer than 10 minutes with under 30 shots. In Appendix C, we present a random sample from the VISTA dataset.
@@ 4 Benchmarking VISTA
A challenge in video-to-text summarization is structuring the generated summaries in a coherent and faithful manner. Directly learning the mapping from $v$ to $s$ could lead to inadequate outputs, as the model lacks explicit guidance on how to organize and present the extracted information. Scientific abstracts often follow a relatively well-defined structure, making them suitable for a more structured generation approach Takeshita et al. (2024). We follow previous work Narayan et al. (2021); Liu et al. (2023); Narayan et al. (2023) in adopting a plan-based framework that introduces an intermediate representation to capture latent structure more effectively than simpler end-to-end approaches. Specifically, given input video $v$, we first generate plan $p$, which consists of a sequence of automatically generated questions $\{q_{1},q_{2},\ldots,q_{m}\}$, each corresponding to a sentence to be verbalized in the summary. The plan explicitly controls the structure of the summary as a whole and the content of each of its sentences (which are meant to answer the questions in the plan). The model is then trained to learn the extended conditional probability distribution $P(s\leavevmode\nobreak\ |\leavevmode\nobreak\ v,p)$, ensuring that the generated summaries follow the structure and flow of plan $p$.
@@
We hypothesize that summary sentences can be viewed as responses to plan questions directly associated with them. This idea is inspired by the theory of Questions Under Discussion (QUD) Roberts (2012); Wu et al. (2023b); Suvarna et al. (2024), which posits that discourse often revolves around a set of questions that guide the structure and interpretation of the conversation.
@@ 6 Results and Analysis
We analyze the plan generation ablation strategy by comparing it with simpler baselines: Lead-3Q, Tail-3Q, and Random-3Q. In these ablation baselines, plans are generated by selecting the first three, last three, or three randomly chosen summary sentences, respectively. Each selected sentence serves as a target for generating a question, with its preceding sentences providing the context. For instance, in the Lead-3Q setting, the first sentence is used as the target (without any preceding context), prompting the first question in the plan, while subsequent sentences incorporate earlier ones as context. Additionally, we compare the case where QUD is not considered. That is, we directly let GPT-o1 generate all planning questions at once only based on the reference summary (NoQUD).
@@
Table 4 underlines the performance differences across different plan generation ablation strategies. NoQUD is also a plan-based approach. It has lower data processing overhead than our original method and performs better than the end-to-end method. However, it still falls short to some extent compared to our approach. The Lead-3Q strategy performs better overall compared to Tail-3Q and Random-3Q, indicating that initial sentences offer stronger contextual continuity for generating plan questions. Nonetheless, these heuristic strategies fail to match the performance of the original planning method.
@@ 8 Conclusion
This paper introduces VISTA, a dataset for summarizing scientific video presentations into concise textual summaries. Comprehensive evaluations across multiple large models demonstrate that the summarization task is challenging, relying on the interplay of multiple modalities (video, text, and audio). We further introduce a plan-based approach, which yields improvements in summary quality and factual accuracy. Beyond dataset creation, our work also confirms that current leading large models still exhibit a noticeable gap compared to human performance.
@@
We use several state-of-the-art large models in our experiments and select the best-performing model, mPLUG-Owl3, to demonstrate the effectiveness of the planning strategy. These large models may carry biases introduced during pretraining. We have not assessed the extent of these biases, as they lie beyond the scope of this study. Furthermore, we have not tested the plan-based approach on all model variants presented in our experiments (e.g., text-based large models and audio-based large models). Our work does not aim to prove that the plan-based method is effective in all models of different modalities, but rather to demonstrate that the plan-based method can improve the performance of video-based models on our dataset. Moreover, plan-based methods can take many different forms, and our work does not aim to identify the optimal planning approach for our dataset. Future work could examine how the plan-based method performs across a wider range of models and modalities.
@@ Acknowledgements
This project has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 Research and Innovation Programme (Grant Agreement No. 948878). We are grateful to the anonymous reviewers and area chairs for their exceptionally detailed and helpful feedback.
@@ Appendix C Data Sample
The VISTA dataset contains carefully curated video-text pairs, predominantly sourced from published papers, aiming to ensure a high standard of quality and relevance. The accompanying texts are designed to function as summaries of their respective videos, offering a concise representation of their content (see Figure 7). Additionally, our dataset focuses on topics within the field of artificial intelligence, making it a good resource for research in AI-related video-to-text summarization and comprehension.